OpenStudy (loser66):

what is the inverse of \[A =\left[\begin{matrix}-1&e^{2t}&0\\1&e^{2t}&0\\0&0&e^{3t}\end{matrix}\right]\] Please, step by step

OpenStudy (loser66):

det A = -2 \(e^{5t}\) and cofactor of it is \[\left[\begin{matrix}e^{5t}&-e^{3t}&0\\-e^{5t}&-e^{3t}&0\\0&0&-2e^{2t}\end{matrix}\right]\]

OpenStudy (loser66):

and then 1/det A *cofactor^(transpose) is \[\frac{-1}{2e^{5t}}\left[\begin{matrix}e^{5t}&-e^{5t}&0\\-e^{3t}&-e^{3t}&0\\0&0&-2e^{2t}\end{matrix}\right]\] \[=\left[\begin{matrix}\frac{-1}{2}&\frac{1}{2}&0\\\frac{1}{2}e^{-2t}&\frac{1}{2}e^{-2t}&0\\0&0&e^{-3t}\end{matrix}\right]\]

OpenStudy (loser66):

so, the last one is \(A^{-1}\), right? and to check whether it's right or wrong, \(AA^{-1}\)should be \(I\), right?

OpenStudy (loser66):

@e.mccormick please help me check my stuff. when checking, I don't get identity matrix

OpenStudy (e.mccormick):

FYI: Transpose of the cofactor is called the adjoint.

OpenStudy (loser66):

FYI??

OpenStudy (e.mccormick):

For Your Information

OpenStudy (loser66):

ok,

OpenStudy (loser66):

@ybarrap

OpenStudy (ybarrap):

Do you know Gauss-Jordan method for finding inverse?

OpenStudy (loser66):

I applied cofactor

OpenStudy (ybarrap):

Can you find inverse any way you want, or do you need to use cofactor technique?

OpenStudy (loser66):

any way If it helps, however, in some case I need cofactor because det A =0, right?

OpenStudy (ybarrap):

If you find detA=0, then there is no inverse.

OpenStudy (ybarrap):

i'm checking result now

OpenStudy (ybarrap):

Your inverse is correct. I also validated to get Identify.

OpenStudy (ybarrap):

I'm going to show you how did with Gauss-Jordan technique: